Two blocks are suspended over a pulley by a string of negligible mass as shown below. The block on the left has a mass of m1, and the block on the right has mass mi, The pulley is a unifonn solid cylinder with mass Mand radius R. The block on the right has a downward acceleration equal to 1/3 the acceleration due to gravity. The tension in the string supporting the mass on the left is T, = 170N and the tension in the string supporting the mass on the right is Ti= 255N. The string does not slip on the pulley.

Two blocks are suspended over a pulley by a string
Answer #1

given

T1 = 170 N

T2 = 255 N

a = g/3 = 9.8/3 = 3.27 m/s^2

11) b. m2 = 39 kg

T2 = m2*g - m2*a

T2 = m2*(g - a)

m2 = T2/(g - a)

= 255/(9.8 - 9.8/3)

= 39 kg

12) e. 52 kg

net Torque = I*alfa

(T2 - T1)*R = 0.5*M*R^2*a/R

T2 - T1 = 0.5*M*a

==> M = 2*(T2-T1)/a

=2*(255 - 170)/(9.8/3)

= 52 kg